# Perturbed Decision-Focused Learning for Modeling Strategic Energy Storage

### v1

@@ I Introduction
Over the past decade, energy storage integration has proven essential for economic and reliable power system decarbonization [1]. However, integrating storage presents unique challenges: energy storage must strategically plan its operations based on future price expectations to quantify its opportunity value, given its limited energy capacity [2, 3]. For instance, in price arbitrage—the most popular grid service for energy storage in North America—storage must accurately predict future prices, especially price spikes, and valleys, to determine the optimal times to charge and discharge. In practice, most energy storage owners adopt a model predictive control (MPC) framework, which decouples price prediction from storage control. They develop proprietary electricity price predictors to capture market fluctuations and input the predicted prices into arbitrage optimization models.
However, decoupling prediction and optimization in MPC frameworks introduces storage control and regulation challenges. For storage operators, machine learning predictors often struggle to effectively account for how their predictions will be utilized in the physical energy storage model, particularly given the limited data availability and computing power in practical implementations [4]. Conversely, electricity market regulators are concerned that non-transparent storage arbitrage models may enable storage operators to exercise market power [5], thereby impeding the convergence of social welfare. Regulators seek systematic approaches to navigate storage operators’ proprietary prediction models and understand the correlation between strategic storage behaviors and market price dynamics.
This paper proposes a decision-focused framework incorporating an energy storage model as a perturbed differentiable layer in the learning pipeline. Our proposed framework offers a significant advantage over existing decision-focused approaches [6] in that it does not rely on ground truth prediction information. The result learning model is more efficient at both predicting and optimizing storage operation from the perspective of storage operators or regulators. The contributions of this paper include:
@@ II-A Learning-aided Storage Operation
Storage owners can arbitrage electricity prices by charging energy when prices are low and selling it back when prices are high. They can submit bids in the day-ahead market (DAM) and the real-time market (RTM), each with its own market clearing timelines and rules. Various methods have been proposed, such as bi-level optimization [7] and dynamic programming [8] [9]. For small-scale energy storage, such as Behind-the-Meter (BTM) energy storage, owners can self-schedule operations without informing the system operator. Many approaches have been proposed, including dynamical programming [10], model predictive control (MPC) [11][12], and reinforcement learning [13]. Electricity price fluctuations are the most critical factor in deciding the best bid or schedule options. Various models have been developed over the years to address this uncertainty, employing techniques such as Markov process models [14], frequency analysis [15], and deep neural networks [16, 17].
@@ II-B Learning-aided Storage Monitoring
The most relevant research is the energy storage model identification [22] [23], which aims to identify the unknown parameters of the energy storage model and uses a pre-built price predictor to forecast the energy storage behavior based on the identified model. On the one hand, knowing the actual price predictor the energy storage owner uses is difficult. Instead, the system operator usually can only observe the historical energy storage behaviors. On the other hand, the price predictor mainly focuses on predicting the mean value of the price trend but falls short of capturing the price variations. The behaviors of energy storage arbitrage are based on the price spread instead of the true price value. Therefore, simply feeding the price prediction into the optimization model may not predict the energy storage behavior well.
@@ III-A Energy Storage Arbitrage Model
where $p_{t}$ and $b_{t}$ are discharge and charge power at time $t$, while $\bar{p}_{t}$ and $\bar{b}_{t}$ are the optimized storage dispatch decisions. $T$ is the time horizon. $\lambda_{t}$ is the predicted real-time price. $e_{t}$ is the state-of-charge (SoC). $P$ is the power rating of storage. $\eta$ is the storage efficiency. $E$ is the SoC capacity. $u(p_{t},b_{t})$ is the charge/discharge cost term and it is a convex function over $p_{t},b_{t}$. For example, $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}+C_{3}b_{t}+C_{4}b_{t}^{2}$ models the linear and quadratic discharging cost. The objective of the arbitrage model is to maximize its profit while keeping the charge/discharge decision within the physical constraints. (1b) sets the upper and lower bounds for the charge and discharge power. (1c) is a sufficient condition to enforce the storage not to charge and discharge at the same time [9]. (1d) sets the upper and lower bounds for SoC. (1e) models the SoC evolution process. In the following, we denote the constraints in (1b)-(1e) as the feasibility set $\mathcal{X}$.
@@ IV-B The Learning Approach
where $\bar{\bm{y}}$ represents the set of target charge and discharge decisions $y_{t},t=1,2,...,T$. It is clear that $\mathcal{L}^{FY}(\hat{\bm{\lambda}},\bar{\bm{y}})\geq 0$ as $\bm{y}^{*}(\hat{\bm{\lambda}})$ is the maximizer given $\hat{\bm{\lambda}}$. The equality $\mathcal{L}^{FY}(\hat{\bm{\lambda}},\bar{\bm{y}})=0$ holds only when $\bm{y}^{*}(\hat{\bm{\lambda}})=\bar{\bm{y}}$. Unfortunately, the loss function in (9) is not continuous because $\bm{y}^{*}(\hat{\bm{\lambda}})$ may not be continuous in $\hat{\bm{\lambda}}$[24][27]. The discontinuity poses computation difficulties when applying backpropagation to update the weights. We employ the perturbation idea in reference [27] by perturbing the reward $\hat{\bm{\lambda}}$ with additive noise. Then the perturbed optimizer is:
@@ IV-C Convex and Smooth Perturbed Loss Function
We will show in the following Propositions that the perturbed loss function is both convex and smooth, which would facilitate efficient training in the learning pipeline. Our work is motivated by recent work [27]. Due to the space limit, we provide detailed proofs to all propositions in the Appendix.
@@
where $\propto$ denotes “proportional to.” Proposition 1 proves that $F_{\epsilon}(\hat{\bm{\lambda}})$ is differentiable and Lipschitz continuous, we demonstrate that by adding appropriate Gaussian noise $\bm{Z}$, the perturbed function becomes differentiable with respect to $\hat{\bm{\lambda}}$. The equations from the first statement of Proposition 1 will be used in subsequent proofs. The Lipschitz continuity of $F_{\epsilon}(\hat{\bm{\lambda}})$ helps establish the Lipschitz continuity of ${\nabla\mathcal{L}_{\epsilon}^{PFY}(\hat{\bm{\lambda}},\bar{\bm{y}})}$ in Proposition 4. Given the perturbation, we aim to understand the impact of $\epsilon$ and how the optimal solution $\bm{y}_{\epsilon}^{*}(\hat{\bm{\lambda}})$ behaves as $\epsilon$ varies. Consequently, we introduce Proposition 2:
@@ IV-D Algorithmic Implementation
Since the exact computation of the expectation in (11) is generally intractable, we approximate it using Monte Carlo sampling in the solution algorithm implementation. The perturbed loss function is approximated as:
@@
where the term $\frac{\partial\hat{\bm{\lambda}}}{\partial\bm{w}}$ represents the Jacobian of reward prediction with respect to the weights $\bm{w}$. The Jacobian can be computed using automatic differentiation [28] in a standard deep learning framework. The details of behavior prediction are outlined in Algorithm 1. The energy storage arbitrage follows a similar process as described in Algorithm 2. Due to space limitations, Algorithm 2 is moved to the Appendix. Readers can refer to the Appendix for more details.
@@ V Experiments
The experiments are implemented in Python 3, Pytorch [29] and PyEPO package [30] on a desktop with 3.0 GHz Intel Core i9, NVIDIA 4080 16 GB Graphic card, and 32 GB RAM. The optimization layer is implemented using the Groubi optimization package [31]. The pre-built predictor is implemented in Keras [32] and Tensorflow [33].
@@ V-A Self-Scheduling Energy Storage Arbitrage
We conduct experiments focusing on Behind-the-Meter energy storage arbitrage, where the energy storage system autonomously manages its scheduling based on future price predictions at hourly real-time prices [12]. The reward forecasting horizon spans 24 hours and is updated hourly. Therefore, the corresponding energy storage schedule is dynamically adjusted on an hourly basis. We use price data from the New York Independent System Operator (NYISO), using data from 2017-2020 for training and 2021 data for testing. The arbitrage profit is computed as
@@
Figure 2: Comparison of annual accumulative profits between the proposed approach and two benchmark methods. The energy storage model is with (a)linear cost term, (b)linear and quadratic cost terms.
We employ the Multi-Layer Perceptron (MLP) as our prediction layer. We compare the arbitrage performance of the proposed method with two MPC-based benchmark methods. The first benchmark employs a Long Short-Term Memory (LSTM) network for real-time price prediction, while the second employs an MLP network. For brevity, we denote these two methods as “LSTM-MPC” and “MLP-MPC”, respectively. We test two types of energy storage models to evaluate the performance of our proposed method: one with a linear cost term and the other with both linear and quadratic cost terms. The linear cost term is $u(p_{t},b_{t})=C_{1}p_{t}$, where the coefficient of discharge cost is $C_{1}=10$. The linear and quadratic cost terms are $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}$, where the coefficients of discharge cost are $C_{1}=5$ and $C_{2}=5$.
@@
Fig. 2 shows the arbitrage performance of the proposed method and the two benchmark methods with the two considered cost terms. The mean absolute error (MAE) of the MLP model is 10.83, while the MAE of the LSTM model is 9.31. Although the lower MAE of the LSTM indicates better overall prediction accuracy compared to the MLP, the MLP model performs better in terms of profit. This is because the MLP predicts the price peak timing more accurately, a crucial factor for optimizing energy storage arbitrage. The proposed method consistently outperforms the two benchmark methods. Our proposed method demonstrates a performance improvement of approximately $47\%$ and $77\%$ over the two benchmark methods with linear cost and 25% and 39% with linear and quadratic cost. Our end-to-end pipeline, which incorporates the optimization model and focuses on decision error, contributes to its profit improvement. The results in the two sub-figures also demonstrate that our proposed method is robust across different energy storage cost models.
@@
Proposed 8809 6381 4589 2990
MLP-MPC 8430 5545 3121 1332
LSTM-MPC 7248 4554 2583 1233

### v2

@@ I Introduction
Over the past decade, energy storage integration has proven essential for economical and reliable power system decarbonization [1]. Facilitated by Federal Energy Regulatory Commission (FERC) Order 841 [2], energy storage systems can now participate in all wholesale energy markets in the United States. Given their limited capacity, energy storage resources must strategically plan their responses based on future price expectations [3], where this process is known as energy storage arbitrage. Price arbitrage has become one of the most popular services [4], whereby energy storage systems discharge when prices are high and charge when prices are low. The energy storage arbitrage provides flexible generation and demand capacity, improving power system efficiency and supporting a more resilient grid.
In practice, most energy storage owners adopt a model predictive control (MPC) framework for storage arbitrage, which decouples price prediction from storage control. They develop proprietary electricity price predictors to capture market fluctuations and input the predicted prices into arbitrage optimization models. However, decoupling prediction and optimization in MPC frameworks introduces storage control and regulation challenges. For storage operators, machine learning predictors often struggle to effectively account for how their predictions will be utilized in the physical energy storage model, particularly given the limited data availability and computing power in practical implementations [5].
Moreover, as each energy storage owner develops their own price predictor to maximize profits, energy storage behaviors are intertwined with electricity pricing dynamics. Given the volatile nature of electricity prices, these arbitrage behaviors exhibit inherent variability, adding uncertainty to the power system and posing significant challenges for grid operation and planning. While energy storage offers promising benefits, its unpredictable actions can disrupt grid stability. Electricity market regulators are increasingly concerned that non-transparent storage arbitrage models may enable operators to exercise market power [6], potentially undermining social welfare. As a result, regulators seek systematic approaches to interpret proprietary prediction models and understand the relationship between strategic storage behaviors and market prices. Accurate prediction of energy storage behavior is thus crucial for power system planning and operations, including net-load forecasting, demand response estimation, market power monitoring, and resource adequacy planning.
To address these challenges, this paper proposes a decision-focused framework incorporating the energy storage behavior as a perturbed differentiable layer in the learning pipeline. Our proposed framework offers a significant advantage over existing decision-focused approaches [7] in that it does not rely on ground truth prediction information. The result learning model is more efficient at both predicting and optimizing storage operation from the perspective of storage operators or regulators. The proposed framework is validated in two strategic applications: (1) energy storage arbitrage, which maximizes profit through optimized storage operations, and (2) energy storage behavior prediction, which forecasts storage actions based on historical data. The contributions of this paper include:
@@ II-A Learning-aided Storage Operation
Storage owners can arbitrage electricity prices by charging energy when prices are low and selling it back when prices are high. They can submit bids in the day-ahead market (DAM) and the real-time market (RTM), each with its own market clearing timelines and rules. Various methods have been proposed, such as bi-level optimization [8] and dynamic programming [9] [10]. For small-scale energy storage, such as Behind-the-Meter (BTM) energy storage, owners can self-schedule operations without informing the system operator. Many approaches have been proposed, including dynamical programming [11], model predictive control (MPC) [12][13], and reinforcement learning (RL) [14]. Electricity price fluctuations are the most critical factor in deciding the best bid or schedule options. Various models have been developed over the years to address this uncertainty, employing techniques such as Markov process models [15], frequency analysis [16], and deep neural networks [17, 18]. In dynamic programming or MPC setups, the prediction model is trained to minimize forecast error, not decision error, resulting in suboptimal profit maximization. Our proposed decision-focused framework incorporates energy storage decision-making into the model training (i.e., training to minimize decision loss) and enhances arbitrage performance. Additionally, our method avoids the inefficiencies of reinforcement learning (RL), RL relies on trial-and-error exploration to optimize policies, which can be time-consuming and inefficient for tasks with well-defined objectives, such as energy storage arbitrage. Our method directly measures how well predictions support optimal decisions, providing a more effective solution. Furthermore, RL cannot handle optimization constraints effectively. RL requires discretizing the action space, and this discretization leads to inefficiencies in representing system dynamics, such as state-of-charge (SoC) transitions, which increase training complexity and reduce interpretability. In contrast, our approach explicitly integrates physical constraints into the model, enabling optimized decision-making and enhancing interpretability.
@@ II-B Learning-aided Storage Monitoring
The most relevant research is the energy storage model identification [23] [24], which aims to identify the unknown parameters of the energy storage model and uses a pre-built price predictor to forecast the energy storage behavior based on the identified model. On the one hand, knowing the actual price predictor the energy storage owner uses is difficult. Instead, the system operator usually can only observe the historical energy storage behaviors. On the other hand, the price predictor mainly focuses on predicting the mean value of the price trend but falls short of capturing the price variations. The behaviors of energy storage arbitrage are based on the price spread instead of the true price value. Therefore, simply feeding the price prediction into the optimization model may not predict the energy storage behavior well. Compared to the energy storage model identification approaches in [23] and [24], our method can infer dynamics of price forecasting directly from historical energy storage decisions. It then uses the predicted reward to estimate future energy storage behavior. Alternatively, one could directly build a neural network to map input features to observed energy storage behaviors. However, such approaches fail to account for the optimization structure of energy storage, leading to degraded prediction performance.
@@ III-A Energy Storage Arbitrage Model
where $p_{t}$ and $b_{t}$ are discharge and charge power at time $t$, while $\bar{p}_{t}$ and $\bar{b}_{t}$ are the optimized storage dispatch decisions. $T$ is the time horizon. $\lambda_{t}$ is the predicted real-time price. $e_{t}$ is the state-of-charge (SoC). $P$ is the power rating of storage. $\eta$ is the storage efficiency. $E$ is the SoC capacity. $u(p_{t},b_{t})$ is the charge/discharge cost term and it is a convex function over $p_{t},b_{t}$. For example, $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}+C_{3}b_{t}+C_{4}b_{t}^{2}$ models the linear and quadratic discharging cost. The objective of the arbitrage model is to maximize its profit while keeping the charge/discharge decision within the physical constraints. (1b) sets the upper and lower bounds for the charge and discharge power. As noted in [29], constraint (1c) is a convex relaxation of the non-convex constraint $p_{t}b_{t}=0$ that prevents simultaneous charging and discharging in energy storage. The non-convex constraint can also be convexified with a small error, as shown in [30]. (1d) sets the upper and lower bounds for SoC. (1e) models the SoC evolution process. In the following, we denote the constraints in (1b)-(1e) as the feasibility set $\mathcal{X}$.
@@ IV-B The Learning Approach
where $\bar{\bm{y}}$ represents the set of target charge and discharge decisions $y_{t},t=1,2,...,T$. $\hat{\bm{\lambda}}^{T}$ denotes the transpose of $\hat{\bm{\lambda}}$. It is clear that $\mathcal{L}^{FY}(\hat{\bm{\lambda}},\bar{\bm{y}})\geq 0$ as $\bm{y}^{*}(\hat{\bm{\lambda}})$ is the maximizer given $\hat{\bm{\lambda}}$. The equality $\mathcal{L}^{FY}(\hat{\bm{\lambda}},\bar{\bm{y}})=0$ holds only when $\bm{y}^{*}(\hat{\bm{\lambda}})=\bar{\bm{y}}$. Unfortunately, the loss function in (9) is not continuous because $\bm{y}^{*}(\hat{\bm{\lambda}})$ may not be continuous in $\hat{\bm{\lambda}}$[25][28]. The discontinuity poses computation difficulties when applying backpropagation to update the weights. We employ the perturbation idea in reference [28] by perturbing the reward $\hat{\bm{\lambda}}$ with additive noise. Then the perturbed optimizer is:
@@ IV-C Convex and Smooth Perturbed Loss Function
We will show in the following Propositions that the perturbed loss function is both convex and smooth, which would facilitate efficient training in the learning pipeline. Our work is motivated by recent work [28]. Due to the space limit, we provide detailed proofs to all propositions in the Appendix[31] [^1] The Appendix can be found at: https://arxiv.org/pdf/2406.17085.
@@
where $\propto$ denotes “proportional to.” Proposition 1 proves that $F_{\epsilon}(\hat{\bm{\lambda}})$ is differentiable and Lipschitz continuous, we demonstrate that by adding appropriate Gaussian noise $\bm{Z}$, the perturbed function becomes differentiable with respect to $\hat{\bm{\lambda}}$. The non-differentiability issue in the non-perturbed problem (4) arises from the linear energy storage model, where linear programming often involves a piecewise linear structure. The solution from the linear energy storage model typically lies on the vertices of a polytope. Consequently, small changes in the input price predictions can cause abrupt shifts in the solutions, leading to non-differentiability of the corresponding loss function. Adding perturbation smooths these abrupt transitions and ensures the perturbed optimization function $F_{\epsilon}(\hat{\bm{\lambda}})$ differentiable. The equations from the first statement of Proposition 1 will be used in subsequent proofs. The Lipschitz continuity of $F_{\epsilon}(\hat{\bm{\lambda}})$ helps establish the Lipschitz continuity of ${\nabla\mathcal{L}_{\epsilon}^{PFY}(\hat{\bm{\lambda}},\bar{\bm{y}})}$ in Proposition 4. The Lipschitz continuity ensures that the function does not change abruptly and the function is continuous. Given the perturbation, we aim to understand the impact of $\epsilon$ and how the optimal solution $\bm{y}_{\epsilon}^{*}(\hat{\bm{\lambda}})$ behaves as $\epsilon$ varies. Consequently, we introduce Proposition 2:
@@ IV-D Algorithmic Implementation
Since the exact computation of the expectation in (11) is generally intractable, we approximate it using Monte Carlo sampling in the solution algorithm implementation. Monte Carlo sampling approximates expectations by averaging over $K$ number of random samples. The perturbed loss function is approximated as:
@@
where the term $\frac{\partial\hat{\bm{\lambda}}}{\partial\bm{w}}$ represents the Jacobian of reward prediction with respect to the weights $\bm{w}$. The Jacobian can be computed using automatic differentiation [32] in a standard deep learning framework. The details of behavior prediction are outlined in Algorithm 1. The energy storage arbitrage follows a similar process as described in Algorithm 2. Due to space limitations, Algorithm 2 is moved to the Appendix [31]. Readers can refer to the Appendix [31] for more details.
@@ V Experiments
The experiments are implemented in Python 3, Pytorch [33] and PyEPO package [34] on a desktop with 3.0 GHz Intel Core i9, NVIDIA 4080 16 GB Graphic card, and 32 GB RAM. The optimization layer is implemented using the Groubi optimization package [35]. The pre-built predictor is implemented in Keras [36] and Tensorflow [37]. The computational time for reward prediction and optimization is less than 0.1 seconds.
@@
In this paper, we employ two neural network models: a Long Short-Term Memory (LSTM) network and a Multi-Layer Perceptron (MLP) network. The LSTM model architecture includes a single LSTM layer, followed by two fully connected layers with a hidden dimension of 64. A ReLU activation function is applied after each fully connected layer. For the MLP model, the architecture consists of three fully connected layers, each with a hidden dimension of 96, with ReLU as the activation function. The Adam optimizer is used for model training. The learning rate is set as 0.01. The input features, such as demand and price data, are structured as time-series historical data, enabling the model to effectively capture temporal dependencies. The neural network outputs a time series of reward predictions, while the optimization layer generates energy storage decisions that align with the reward prediction horizon. In our experiments, we used a 24-hour look-back window as input, and the model predicts the reward and corresponding storage decisions for the subsequent 24-hour period.
@@ V-A Self-Scheduling Energy Storage Arbitrage
We conduct experiments focusing on Behind-the-Meter energy storage arbitrage, where the energy storage system autonomously manages its scheduling based on future price predictions at hourly real-time prices [13]. The reward forecasting horizon spans 24 hours and is updated hourly. Therefore, the corresponding energy storage schedule is dynamically adjusted on an hourly basis. We collect data from the New York Independent System Operator (NYISO), using data from 2017-2020 for training and 2021 for testing. The historical data for Real-Time Price (RTP), Day-Ahead Price (DAP), and load data are collected. The data resolution is one sample per hour. The dataset includes both input features and target decision values. The target decisions are obtained by solving (1) using ground truth RTP. The input features consist of three components: RTP, DAP, and load data. These input features are typical for training a price forecaster. To augment the training samples in the dataset, a rolling horizon window with a step size of one hour is employed. The input features of the training dataset are structured as a tensor of size $35017\times 24\times 3$, where 35017 represents the number of samples, 24 corresponds to hourly intervals, and 3 denotes the number of feature types. The target decision data is represented as a $35017\times 24$ matrix. The testing dataset follows the same format, comprising 8713 samples.
The arbitrage profit is computed as
@@
Figure 2: Comparison of annual accumulative profits between the proposed approach and three benchmark methods. The energy storage model is with (a) linear cost term, (b) linear and quadratic cost terms.
The goal of this experiment is to compare the proposed decision-focused approach with several data-driven benchmark approaches to demonstrate that training with our decision-focused loss leads to increased profit. We compare the cumulative profit over the entire year of 2021 and show the performance improvements achieved by our method. We employ the MLP as our prediction layer. We compare the arbitrage performance of the proposed method with two MPC-based benchmark methods. The first benchmark employs an LSTM network for real-time price prediction, while the second employs an MLP network. For brevity, we denote these two methods as “LSTM-MPC” and “MLP-MPC”, respectively. We also include a benchmark method based on reinforcement learning (RL) [14], which employs a Markov decision process (MDP) framework by discretizing the SoC. The original RL approach assumes perfect energy storage efficiency ($\eta=1$). To ensure a fair comparison, we revised the RL approach by incorporating the SoC transition into the state transition of the RL framework and set the storage efficiency of all methods to 0.9. We test two types of energy storage models to evaluate the performance of our proposed method: one with a linear cost term and the other with both linear and quadratic cost terms. The linear cost term is $u(p_{t},b_{t})=C_{1}p_{t}$, where the coefficient of discharge cost is $C_{1}=10$. The linear and quadratic cost terms are $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}$, where the coefficients of discharge cost are $C_{1}=5$ and $C_{2}=5$.
@@
Fig. 2 shows the arbitrage performance of the proposed method and the three benchmark methods with the two considered cost terms. The mean absolute error (MAE) of the MLP model is 10.83, while the MAE of the LSTM model is 9.31. Although the lower MAE of the LSTM indicates better overall prediction accuracy compared to the MLP, the MLP model performs better in terms of profit. This is because the MLP predicts the price peak timing more accurately, a crucial factor for optimizing energy storage arbitrage. The proposed method consistently outperforms the three benchmark methods. Our proposed method demonstrates a performance improvement of approximately $47\%$, $71\%$, and $50\%$ over the three benchmark methods with linear cost and 30%, 38%, and 72% with linear and quadratic cost. Our end-to-end pipeline, which incorporates the optimization model and focuses on decision error, contributes to its profit improvement. The results in the two sub-figures also demonstrate that our proposed method is robust across different energy storage cost models.
@@
Proposed 9126 6381 4589 2990
MLP-MPC 8388 5545 3121 1332
LSTM-MPC 6910 4554 2583 1233
In Fig. 3, we present the predictions of the proposed method, LSTM, and MLP models, and the ground truth real-time price (RTP) over two consecutive days. We also show the corresponding energy storage decisions based on these predictions. As shown in Fig. 3, while the reward predictions from all three methods exhibit similar trends, our method provides more accurate timing of price spikes, resulting in the best decision-making outcomes. The energy storage behavior from our method aligns closely with the optimal decisions. In contrast, the decisions generated by the LSTM and MLP models deviate from the optimal decisions and show larger errors. The profits for this segment are \$$\$363$, \$$\$220$, and \$$\$204$ for our method, MLP method, and LSTM method, respectively.
Figure 3: Comparison of reward predictions and corresponding decisions between the proposed approach, two benchmark methods, and ground truth. The first subfigure shows the reward predictions and the ground truth real-time price (RTP). The subsequent subfigures show the corresponding decisions compared to the optimal decisions.
To assess the adaptability of the proposed method to a new energy storage system, we employ an energy storage system with parameters $\eta=0.95$, $E=1$, $P=0.5$, $C_{1}=10$, and $C_{2}=0$. A reward predictor is trained using the optimal decisions generated by this energy storage model. The trained predictor is then applied to a different model with parameters $\eta=0.9$, $E=2$, $P=0.5$, $C_{1}=5$, and $C_{2}=5$. The arbitrage profit achieved through adaptation is \$4,783, compared to \$5,314 from the original model. We abbreviate the proposed method, trained on a storage system with different parameters, as “Proposed-Adapt.” The performance comparison in Fig. 4 shows that the algorithm adapts effectively to the new model, with only a minor reduction in profit. This demonstrates the robustness of the trained predictor in handling variations in system parameters while maintaining competitive performance.
Figure 4: Arbitrage performance based on training on the original energy storage system and on a storage system with different parameters.